OpenStudy (itsmichelle29):

help plz medal and fan Simplify sin^2 theta /1- cos^2 theta

Nnesha (nnesha):

\[\huge\rm cos^2 \theta + \sin^2 \theta =1 \]you should remember this equation!!

Nnesha (nnesha):

solve for sin^2theta

OpenStudy (anonymous):

-cos(2 theta) 2 sin^2(theta)-1

OpenStudy (itsmichelle29):

cos^2 theta .. yes ????

Nnesha (nnesha):

what about it ? :=)

OpenStudy (itsmichelle29):

is that the answer ????

Nnesha (nnesha):

no

OpenStudy (itsmichelle29):

Okay let me try and get it

Nnesha (nnesha):

sure!

imqwerty (imqwerty):

sin^2theta +cos^2theta = 1 1-cos^2theta=sin^2theta so we can write -> sin^2theta/sin^2theta = 1 but the condition is that sin(theta) does not equal 0

OpenStudy (itsmichelle29):

sin2theta

OpenStudy (itsmichelle29):

Yes ??

Nnesha (nnesha):

that's one of the option....

imqwerty (imqwerty):

can u explain how did u get sin2theta

OpenStudy (itsmichelle29):

By 1-cos2theta=sin2theta........................... i think

Nnesha (nnesha):

\[\frac{ \sin^2\theta }{\color{ReD}{ 1-\cos^2 }}\] now replace 1-cos^2for sin^2x bec 1-cos^2x=sin^2x

OpenStudy (itsmichelle29):

Omg i dont understand

Nnesha (nnesha):

\[\huge\rm cos^2 \theta + \sin^2 \theta =1 \] solve for sin^2\[\rm sin^2\theta=1-\cos^2\theta\] you will get sin^2 =1-cos^2 right ?

OpenStudy (itsmichelle29):

Yes

Nnesha (nnesha):

okay so if sin^2theta =1-cos^2 you can replace the (1-cos^2) which is at the dneominator with sin^2 theta right ?

Nnesha (nnesha):

\[\huge\rm \frac{ \sin^2\theta }{\color{ReD}{ 1-\cos^2 }}\] replace 1-cos^2for sin^2x bec \[\huge\rm \color{reD}{ 1-\cos^2x}=\sin^2x \]

OpenStudy (itsmichelle29):

so the answer is sin right

Nnesha (nnesha):

\[\frac{ \sin^2 x }{ \color{reD}{\sin^2x}}=?\]

Nnesha (nnesha):

let sin^2x = a so a/a = ??

Nnesha (nnesha):

a divide by a = ??

OpenStudy (itsmichelle29):

0

Nnesha (nnesha):

no never!

Nnesha (nnesha):

there is always invisible one!!

Nnesha (nnesha):

2/2 = ??

OpenStudy (itsmichelle29):

1 lol

Nnesha (nnesha):

ye.-.

OpenStudy (itsmichelle29):

okay thnks